What is the probability that the shelf life of the battery is between 400 days and 550 days?


1
Expert's answer
2021-03-09T07:51:57-0500

A certain type of battery has a mean shelf life of 600 days with a standart deviation of 28 days.


The probability that 400<X<550 is equal to the blue area under the curve.



Since "\mu=600"  and "\sigma=28"  we have:

"P(400<X<550) = P(400-600<X-\mu<550-600)\\=P(\frac{400-600}{28}<\frac{X-\mu}{\sigma}<\frac{550-600}{28})"

Since 

"\frac{x-\mu}{\sigma} = Z" , "\frac{400-600}{28} = -7.14" and "\frac{550-600}{28} = -1.79"  we have:

"P(400<X<550)=P(-7.14<Z<-1.79)"

Use the standard normal table to conclude that:


"P(-7.14<Z<-1.79)=P(-1.79)-P(-7.14)=0.03673-0=0.03673"

Answer: 0.0367.